What is the growth factor for the exponential function f(x)=3(1.5)^x
let g(x)=200 (1.2)^x
g(0)=?
g(3)=?

Answer by math_tutor2020(3817)


The growth factor of f(x) = 3(1.5)^x is 1.5
It is the base of the exponential function

In general we have
y = a*b^x
a = initial value = y intercept
b = growth factor if b > 1, decay factor if 0 < b < 1

========================================================

For your other question, we replace x with the values mentioned and compute.
g(x) = 200(1.2)^x
g(0) = 200(1.2)^0
g(0) = 200(1)
g(0) = 200
and
g(x) = 200(1.2)^x
g(3) = 200(1.2)^3
g(3) = 200(1.728)
g(3) = 345.6